### v1

# THEME : Enhancing Thematic Investing with Semantic Stock Representations and Temporal Dynamics

Abstract.
Thematic investing aims to construct portfolios aligned with structural trends, yet selecting relevant stocks remains challenging due to overlapping sector boundaries and evolving market dynamics. To address this challenge, we construct the Thematic Representation Set (TRS), an extended dataset that begins with real-world thematic ETFs and expands upon them by incorporating industry classifications and financial news to overcome their coverage limitations. The final dataset contains both the explicit mapping of themes to their constituent stocks and the rich textual profiles for each. Building on this dataset, we introduce THEME, a hierarchical contrastive learning framework. By representing the textual profiles of themes and stocks as embeddings, THEME first leverages their hierarchical relationship to achieve semantic alignment. Subsequently, it refines these semantic embeddings through a temporal refinement stage that incorporates individual stock returns. The final stock representations are designed for effective retrieval of thematically aligned assets with strong return potential. Empirical results show that THEME outperforms strong baselines across multiple retrieval metrics and significantly improves performance in portfolio construction. By jointly modeling thematic relationships from text and market dynamics from returns, THEME provides a scalable and adaptive solution for navigating complex investment themes.
Hierarchical Contrastive Learning; Thematic Investing; Information Retrieval; Stock Selection; Representation Learning
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 34th ACM International Conference on Information and Knowledge Management; November 10–14, 2025; Seoul, Republic of Korea.booktitle: Proceedings of the 34th ACM International Conference on Information and Knowledge Management (CIKM ’25), November 10–14, 2025, Seoul, Republic of Koreaisbn: 979-8-4007-2040-6/2025/11doi: 10.1145/XXXXXX.XXXXXXccs: Applied computing Economicsccs: Information systems Information retrieval
1. Introduction
Figure 1. Comparison of t-SNE visualizations of stock description embeddings. (Upper Left) A baseline embedding model fails to cluster stocks by thematic meaning. (Lower Right) Our tuned embeddings organize stocks by economic theme, enabling more interpretable and structured retrieval.
Figure 2. Overview of how the THEME system is used in practice and how it works. Given a user-provided thematic description—such as ”AI software and chipmakers”—THEME semantically embeds the input and retrieves stocks that are both thematically aligned and investment-suitable. These selected stocks can then be used to construct a portfolio for real-world investment.
Thematic investing is a popular strategy that constructs portfolios around structural trends such as artificial intelligence, renewable energy, or cybersecurity. Unlike sector-based investing, thematic strategies are inherently cross-sectoral and dynamic: relevant companies often span diverse industries and change over time in response to innovation, regulation, and market sentiment.
Despite its appeal, current thematic investment methods rely heavily on static ETF compositions or expert-defined lists. These approaches lack adaptability and are often slow to reflect emerging companies or shifts in thematic relevance. As a result, investors may miss out on timely opportunities or hold outdated portfolios that no longer represent the target theme.
To address this gap, we propose THEME, a scalable system for thematic stock retrieval and portfolio construction. As illustrated in Figure 1, traditional embedding models trained on general-purpose text fail to produce meaningful clusters of thematically related stocks. This highlights a key limitation: financial texts contain domain-specific semantics that general embeddings cannot fully capture. Our method addresses this by producing domain-tuned embeddings that reflect thematic structure, enabling more interpretable and effective retrieval.
Figure 2 shows how THEME operates in practice. Given a user-defined thematic query, THEME encodes the theme, retrieves semantically and financially relevant stocks, and returns candidates suitable for portfolio construction. The system integrates seamlessly into real-world workflows, supporting both discretionary screening and systematic portfolio updates.
At the core of THEME is a hierarchical contrastive learning framework that combines two sources of supervision. First, semantic signals derived from financial text (e.g., business descriptions, regulatory filings, and news) provide structural thematic alignment. Second, temporal signals based on recent returns guide the model to favor stocks with short-term investment relevance. This two-stage process enables the model to capture both long-term thematic meaning and dynamic market responsiveness.
To support robust learning and generalization, we also construct the Thematic Representation Set (TRS)—a dataset that extends beyond real-world ETF holdings by incorporating sectoral taxonomies and news-derived themes. TRS enables THEME to support hundreds of broad and niche themes, including those not covered by existing ETF products.
In summary, we make the following contributions:
  • A unified semantic-temporal framework for theme modeling: We propose a contrastive learning approach that embeds stocks using both textual descriptions and recent return behavior. This enables theme-aware retrieval that captures both structural meaning and short-term responsiveness.
  • A domain-tuned embedding space for finance: We show that general-purpose embeddings fail to organize stocks by thematic relevance. In contrast, our fine-tuned embeddings yield clearer semantic structure, improving both interpretability and retrieval quality.
  • A hierarchical modeling strategy for thematic structure: Our method supports a wide range of themes, from broad categories to niche subthemes, by applying a multi-level contrastive framework that reflects the layered nature of thematic semantics.
  • A scalable thematic dataset (TRS): We construct the TRS, a large-scale stock-theme dataset that extends beyond ETF coverage by incorporating industry taxonomies and financial news. This improves coverage of emerging and underrepresented themes.
  • Real-world evaluation for portfolio construction: We implement and assess a full-system pipeline that improves retrieval precision and portfolio performance across multiple model backbones. The system is compatible with both discretionary and systematic investment workflows.
2. Related Works
The rise of thematic investing in recent years has prompted research into systematic methods for identifying theme-relevant companies. A key challenge is that themes often cut across conventional industry boundaries. Prior approaches have typically relied on expert curation or static industry classifications, which struggle to capture the full breadth of relevant firms. To address this, researchers have increasingly turned to natural language processing (NLP) techniques applied to financial text.
For instance, Rao et al. (Rao et al., 2023) demonstrate that fit-for-purpose language models can effectively associate companies with long-term themes by comparing thematic keywords to business descriptions. Similarly, Vamvourellis et al. (Vamvourellis et al., 2023) show that embeddings derived from SEC filings using fine-tuned transformer models not only reproduce standard industry groupings but also reveal new cross-sector relationships. These semantic embeddings have enabled applications such as clustering companies by thematic exposure and constructing thematic baskets. Building on this line of work, Takayanagi et al. (Takayanagi et al., 2024) introduce a hybrid approach (SETN) that incorporates both textual and network-based signals to enhance stock representations for thematic investing.
However, financial markets are inherently non-stationary, and a company’s thematic alignment may shift over time due to structural changes, strategic pivots, or macroeconomic developments (Li et al., 2023a). Traditional static models fail to capture this temporal dynamic. In response, recent work has explored time-aware representations. For example, Hwang et al. (Hwang et al., 2023) propose SimStock, a self-supervised framework that generates temporally-sensitive stock embeddings, improving performance on tasks such as thematic index tracking. In practice, dynamic thematic strategies are also gaining popularity, frequently updating exposures based on short-term signals such as news sentiment and price momentum (Gu et al., 2020).
Meanwhile, large language models (LLMs) like GPT-4 have shown growing utility in financial applications (Lee et al., 2023; Kim et al., 2023; Nie et al., 2024; Lee et al., 2025, 2024). In theory, these models can infer thematic relevance from contextual knowledge alone (OpenAI, 2023). Yet, general-purpose LLMs often fall short in this domain: they may lack access to up-to-date disclosures or fail to detect subtle, domain-specific quantitative signals, leading to inconsistent or incomplete outputs (Yang et al., 2023). BloombergGPT (Wu et al., 2023) highlights the potential of domain-specific models, but its application to thematic stock selection remains unexplored.
Despite these advances, no existing system fully integrates semantic NLP techniques with explicit temporal modeling to support dynamic thematic investing (Wang et al., 2023a). This motivates the development of THEME, a unified framework that combines language-based stock embeddings with time-aware performance modeling. By capturing both the textual signals of thematic relevance and their evolution over time, THEME aims to provide a practical and robust solution for identifying and tracking theme-aligned stocks in a non-stationary market environment.
Figure 3. Overview of the two-stage hierarchical contrastive learning framework. By leveraging the explicit relationship between a theme and its known constituent stocks as the primary training signal, the framework first learns semantic embeddings via contrastive learning. Subsequently, a lightweight adapter refines these semantic embeddings by fusing them with temporal embeddings derived from recent stock returns. The resulting model generates theme-aware stock representations that jointly capture semantic meaning and temporal signals, suitable for dynamic portfolio construction.
3. Methodology
This work introduces a hierarchical contrastive learning framework for thematic investing (Figure 3). This framework generates stock embeddings that capture both semantic alignment with theme descriptions and sensitivity to short-term return patterns. Prior contrastive methods often focus solely on static textual similarity. In contrast, our approach jointly models semantic meaning and temporal signals to support dynamic, theme-aware stock retrieval. It is designed to scale to thousands of stocks across asset classes and is optimized for real-world deployment.
We refer to this framework as hierarchical contrastive learning for two key reasons. First, it directly utilizes the hierarchical relationship between themes and their individual constituent stocks in the learning process. Second, it applies two levels of contrastive objectives hierarchically. The first stage is a semantic objective that aligns stocks with the conceptual meaning of a theme, followed sequentially by a temporal objective that refines the embeddings using recent return behavior. This multi-layered hierarchical structure enables the progressive refinement of stock representations from static thematic semantics to dynamic investment relevance.
To support this two-stage learning process, we construct the TRS. TRS begins with 1,153 real-world thematic ETFs, each linked to a textual description and a list of constituent stocks. While ETFs offer strong supervision signals, their coverage is incomplete and biased toward trending sectors such as IT and clean energy. To address this, we incorporate sector and industry classification systems to expand our theme universe to approximately 200 unique themes.
Each TRS record contains a theme label, a textual summary, and its associated list of stocks. For each stock, we construct a rich textual profile by aggregating data from SEC filings and financial news, allowing stocks to participate in multiple themes. This contrasts with traditional systems where each stock belongs to a single sector. Moreover, TRS is updated dynamically with live information, ensuring continued relevance as new themes emerge.
3.1. Stage 1: Thematic Alignment
In the first stage, we learn a shared embedding space that semantically aligns the textual profiles of stocks with the descriptions of themes. This process aims to position semantically related stocks and themes closer together within the embedding space.
First, we use a large-scale, pre-trained text embedding model with frozen parameters as the backbone, denoted by $f_{\Phi}(\cdot)$, where $\Phi$ represents the pre-trained weights. We apply LoRA (Low-Rank Adaptation) (Hu et al., 2022) to this backbone model, introducing a small set of trainable parameters, $\theta$. The final, combined model is denoted as $f^{\prime}_{\theta}(\cdot)$, representing the LoRA-adapted backbone $f_{\Phi}(\cdot)$.
Given a set of theme descriptions $\mathcal{T}=\{t_{1},t_{2},\dots,t_{M}\}$ and a set of stock textual profiles $\mathcal{S}=\{s_{1},s_{2},\dots,s_{N}\}$, the final model $f^{\prime}_{\theta}(\cdot)$ is used to transform each text into a $d$-dimensional vector:
(1) $z_{i}=f^{\prime}_{\theta}(t_{i})\in\mathbb{R}^{d},\quad h_{j}=f^{\prime}_{\theta}(s_{j})\in\mathbb{R}^{d}$
where $z_{i}$ is the embedding for a theme $t_{i}\in\mathcal{T}$, and $h_{j}$ is the embedding for a stock $s_{j}\in\mathcal{S}$.
The learning process uses each theme embedding $z_{i}$ as an anchor to pull closer the embedding of a constituent stock $h_{j}^{+}$ from the theme, which serves as the positive pair, while pushing away the embeddings of non-constituent stocks $h_{j}^{-}$ as negative pairs.
The semantic alignment loss $\mathcal{L}_{\text{align}}$ is given by:
(2) $\mathcal{L}_{\text{align}}=-\log\frac{\exp(\text{sim}(z_{i},h_{j}^{+})/\tau)}{\exp(\text{sim}(z_{i},h_{j}^{+})/\tau)+\sum_{k}\exp(\text{sim}(z_{i},h_{j}^{-})/\tau)}$
Here, sim(,)($\cdot$$\cdot$) is cosine similarity and $\tau$ is the temperature hyperparameter. The efficiency of LoRA enables the rapid and dynamic update of the resulting semantic embeddings to reflect new themes or evolving semantic contexts.
3.2. Stage 2: Temporal Refinement
The second stage refines the semantic embeddings by incorporating dynamic, short-term return signals. To achieve this, we introduce a lightweight 2-layer adapter, denoted as $\mathcal{A}_{\phi}$, with trainable parameters $\phi$. This adapter is designed to fuse semantic context with dynamic temporal patterns. It takes two distinct inputs for a stock $s_{j}$: its semantic embedding ${h}_{j}$ from Stage 1, and its past $L$ trading days of daily returns, ${r}_{j}$, where $L$ is the lookback period. The adapter jointly processes these inputs to produce a single, temporally-refined stock embedding ${h}^{\prime}_{j}$:
(3) ${h}^{\prime}_{j}=\mathcal{A}_{\phi}({h}_{j},{r}_{j})$
The adapter’s parameters $\phi$ are optimized via a triplet loss function designed to rank stocks within a theme based on future returns. For a given theme $t_{i}$, we first select a positive sample $s_{p}$ and a negative sample $s_{n}$ from its constituent stocks. The stock with the higher forward return over a horizon of $H$ days is designated as the positive sample, while the other serves as the negative. These samples, along with the theme’s semantic embedding $z_{i}$ as an anchor, form a training triplet $(z_{i},h^{\prime}_{p},h^{\prime}_{n})$, where $h^{\prime}_{p}$ and $h^{\prime}_{n}$ are the temporally-refined embeddings of the positive and negative stocks, respectively. The training objective is to minimize the following loss, thereby ensuring that the anchor is closer in the embedding space to the positive sample than to the negative sample by at least a margin $m$:
(4) $\mathcal{L}_{\text{triplet}}=\left[\text{sim}\left({z}_{i},{h}^{\prime}_{n}\right)-\text{sim}\left({z}_{i},{h}^{\prime}_{p}\right)+m\right]_{+}$
This approach allows the model to learn a nuanced representation that is not only thematically relevant but also sensitive to recent market dynamics, all while preserving the rich knowledge in the pre-trained semantic embeddings. Additional training details are provided in Section 4.
3.3. Stage 3: Inference Pipeline
At inference time, a user query $q$ is encoded into a query embedding $z_{q}$ via a semantic model $f^{\prime}_{\theta}(\cdot)$. We perform retrieval by ranking a corpus of pre-computed stock embeddings ${h}^{\prime}_{j}$ based on their cosine similarity to the query vector $z_{q}$. The resulting top-$K$ list is subsequently utilized for downstream financial applications, including thematic screening, portfolio construction, and index design.
3.4. System Implementation and Integration
Our system is implemented as a modular, cloud-native application. Embedding and similarity computation scale linearly with the number of stocks, enabling fast inference across global universes. REST APIs expose the core functionality, supporting integration with discretionary research platforms, automated portfolio engines, and personalized investing tools. Feedback mechanisms allow iterative refinement based on user inputs.
The platform is also extensible to real-time signals such as ESG events, earnings announcements, or patent activity. These extensions enable advanced use cases including event-driven rebalancing, theme-based risk monitoring, and semantic pre-filtering for financial NLP pipelines.
Table 1. Comparison of retrieval performance across various methods before and after applying THEME. Metrics include Hit Rate (HR) and Precision (P) at $k\in\{3,5,10\}$, where HR@$k$ denotes the fraction of queries for which at least one relevant item appears in the top-$k$ results, and P@$k$ measures the proportion of relevant items among the top-$k$ retrieved.
Model Size HR@3 P@3 HR@5 P@5 HR@10 P@10
Vanilla Ours Vanilla Ours Vanilla Ours Vanilla Ours Vanilla Ours Vanilla Ours
[…]
bge-small-en-v1.5 (Xiao et al., 2023) 33M 0.1392 0.6031 0.0584 0.3557 0.1701 0.7423 0.0495 0.3722 0.2577 0.8247 0.0541 0.3701
bilingual-embedding-large (Conneau et al., 2019) 559M 0.1598 0.6443 0.0687 0.3935 0.1959 0.7371 0.0691 0.3918 0.3041 0.8093 0.0686 0.3825
multilingual-e5-large-instruct (Wang et al., 2024) 560M 0.1753 0.6959 0.0704 0.4588 0.2577 0.7732 0.0773 0.4392 0.3505 0.8660 0.0789 0.4206
stella_en_1.5B_v5 (Zhang et al., 2025) 1.5B 0.2474 0.7320 0.1375 0.5275 0.3557 0.8299 0.1412 0.5237 0.4536 0.8814 0.1387 0.4856
gte-Qwen2-1.5B-instruct (Li et al., 2023b) 1.5B 0.3144 0.6186 0.1684 0.4124 0.3763 0.6959 0.1608 0.4031 0.4639 0.8093 0.1464 0.3907
SFR-Embedding-Mistral (Meng et al., 2024) 7B 0.3711 0.7887 0.2062 0.6082 0.4639 0.8557 0.2052 0.5845 0.5567 0.9124 0.1995 0.5619
GritLM-7B (Muennighoff et al., 2024) 7B 0.0825 0.8196 0.0344 0.6014 0.1237 0.8814 0.0371 0.5907 0.1753 0.9124 0.0320 0.5701
gte-Qwen2-7B-instruct (Li et al., 2023b) 7B 0.5206 0.7938 0.3299 0.5790 0.6031 0.8351 0.3227 0.5680 0.7165 0.9072 0.2985 0.5366
e5-mistral-7b-instruct (Wang et al., 2023b) 7B 0.3454 0.7887 0.1770 0.5962 0.4691 0.8763 0.1907 0.5794 0.5412 0.9330 0.1912 0.5552
Linq-Embed-Mistral (Choi et al., 2024) 7B 0.5155 0.8196 0.3522 0.6289 0.5773 0.8918 0.3340 0.6041 0.6546 0.9278 0.3155 0.5701
4. Experiments
We design a series of experiments to evaluate the effectiveness of our proposed system, THEME, in both retrieval quality and investment utility. First, we evaluate how accurately the model retrieves stocks relevant to a given theme. This measures the semantic alignment between theme descriptions and candidate stocks, using metrics such as Hit Rate (HR) and Precision (P) at various cutoff thresholds. Second, we assess whether the retrieved stocks lead to improved investment outcomes. We construct portfolios based on the top-ranked results and compare their performance against those generated by baseline methods. Third, we perform an ablation study to analyze the role of theme representation. In particular, we compare two anchor strategies: (i) using a group of representative stocks, and (ii) using a textual theme description.
4.1. Experimental Setup
We train the model using a two-tier dataset. The first component consists of 1,153 real-world thematic ETFs covering approximately 3,000 unique U.S. equities. To ensure quality, we filter for ETFs with at least 10 constituents. The remaining 969 ETFs are randomly split into training (678), validation (97), and test (194) sets. While the validation and test sets are preserved for real-world evaluation, the training data is augmented via our TRS methodology to a final set of 196 themes. This allows the model to learn from a diverse thematic universe while being tested on true market data. For the temporal refinement stage, we use two years of historical U.S. market data, generating training samples via a rolling window with a lookback period of $L=60$ trading days. Pairs of stocks within each theme are labeled based on their relative forward returns over a horizon of $H=14$ days.
4.2. Experiment 1: Retrieval Performance
We compare THEME against multiple baselines including embedding models (e.g., Fin-E5, SFR-Embedding-Mistral), open LLMs (GPT-4.1, Gemini-2.5), and our own Linq-Embed-Mistral. Each model ranks a shared stock universe given a theme description, and we evaluate retrieval quality at $k{=}\{3,5,10\}$ using HR and P.
As shown in Table 1, THEME significantly improves retrieval quality across all model backbones. For example, when applied to Linq-Embed-Mistral, THEME improves HR@3 from 0.5155 to 0.8196 and P@3 from 0.3522 to 0.6289. Similar gains are observed across other 7B-scale models such as SFR and e5-mistral. Even smaller base models (e.g., bge-small-en-v1.5, multilingual-e5) show substantial performance boosts when integrated with THEME, often exceeding the retrieval quality of much larger models in their vanilla form.
Notably, THEME-enhanced versions consistently outperform strong LLM baselines such as GPT-4.1 and Gemini-2.5, which are not directly compatible with contrastive fine-tuning. This highlights the benefit of incorporating theme-aligned supervision at scale. Overall, these results confirm that THEME is effective at improving thematic relevance in stock retrieval tasks, regardless of the underlying model size.
Table 2. Model performance comparison across key financial metrics. Each column reports performance at different top-$k$ cutoffs ($k\in\{3,5,10\}$), with Sharpe Ratio (SR) and Cumulative Return (CR) indicating profitability (higher is better, $\uparrow$), and Maximum Drawdown (MDD) indicating risk (lower is better, $\downarrow$).
Model Type SR@3 SR@5 SR@10 MDD@3 MDD@5 MDD@10 CR@3 CR@5 CR@10
Linq-Embed-Mistral Vanilla 0.4870 0.4530 0.4991 -0.2551 -0.2618 -0.2534 0.0907 0.0831 0.0938
Ours 0.5881 0.5913 0.5432 -0.2526 -0.2474 -0.2440 0.1187 0.1176 0.1043
gte-Qwen2-7B-instruct Vanilla 0.5014 0.4293 0.4712 -0.2427 -0.2413 -0.2392 0.0917 0.0749 0.0843
Ours 0.7592 0.7711 0.6893 -0.2378 -0.2320 -0.2338 0.1645 0.1650 0.1422
GritLM-7B Vanilla 0.5744 0.5912 0.6324 -0.2431 -0.2409 -0.2407 0.1154 0.1169 0.1273
Ours 0.5952 0.5163 0.5291 -0.2683 -0.2589 -0.2463 0.1196 0.0981 0.1007
4.3. Experiment 2: Portfolio Construction
Building on the results from Experiment 4.2, which demonstrated that THEME effectively retrieves thematically relevant stocks, we now assess whether these stocks also yield superior investment performance. To this end, we evaluate the performance of portfolios constructed using the top-$K$ stocks selected by THEME over a rolling test period from April 23, 2024 to April 29, 2025.
For each window, we construct equal-weighted portfolios using the top-$K$ stocks ($K\in\{3,5,10\}$) ranked by similarity to the theme anchor. The average daily return of these $K$ stocks is then recorded over the subsequent 14 trading days. Chaining the returns across all windows produces a continuous daily return series for the entire test period.
Portfolio performance is evaluated with the following metrics:
  • Cumulative Return (CR):
    $\mathrm{CR}_{t}=\prod_{i=1}^{t}(1+r_{i})-1,$
    where $r_{i}$ is the portfolio return on day $i$.
  • Sharpe Ratio (SR):
    $\mathrm{SR}=\frac{\bar{r}}{\sigma_{r}}\cdot\sqrt{252},$
    where $\bar{r}$ and $\sigma_{r}$ are the mean and standard deviation of daily returns.
  • Maximum Drawdown (MDD):
    $\mathrm{MDD}=\min_{t}\left(\frac{V_{t}}{\max_{s\leq t}V_{s}}-1\right),$
    with $V_{t}$ representing cumulative portfolio value at time $t$.
As summarized in Table 2, THEME consistently improves portfolio quality over vanilla approaches. For example, with the gte-Qwen2-7B-instruct model, SR@3 increases from 0.5014 to 0.7592, CR@3 from 0.0917 to 0.1645, and MDD improves from –0.2427 to –0.2378. These results also outperform the baseline performance of investing in the real thematic ETF universe, which achieves a Sharpe ratio of 0.4845, CR of 0.0672, and MDD of –0.2368. Similar advantages are observed across other $K$ values and model variants.
Even with strong baselines such as GritLM-7B, THEME maintains or slightly improves profitability metrics while remaining robust across architectures. These findings indicate that combining semantic and temporal signals enables the retrieval of stocks that are not only thematically aligned but also effective in constructing more profitable and resilient portfolios in dynamic markets.
4.4. Experiment 3: Ablation on Anchor and Dataset Strategy
We conduct an ablation study to investigate how two key aspects of our contrastive learning setup affect retrieval quality: (i) the choice of anchor during training, and (ii) the composition of the training dataset. Each factor is evaluated independently using a consistent set of backbone models and evaluation metrics.
Table 3. Ablation study on anchor selection for semantic alignment. We compare our setting (anchor: theme, positive: stock) with the counterpart (anchor: stock, positive: stock). We report the score improvements of our setting over the counterpart for each evaluation metric.
Model P@3 P@5 P@10
bge-small-en-v1.5 (Xiao et al., 2023) $\uparrow$ 0.1066 $\uparrow$ 0.1433 $\uparrow$ 0.1701
bilingual-embedding-large (Conneau et al., 2019) $\uparrow$ 0.1151 $\uparrow$ 0.1227 $\uparrow$ 0.1243
multilingual-e5-large-instruct (Wang et al., 2024) $\uparrow$ 0.1375 $\uparrow$ 0.1237 $\uparrow$ 0.1361
stella_en_1.5B_v5 (Zhang et al., 2025) $\uparrow$ 0.1083 $\uparrow$ 0.1340 $\uparrow$ 0.1371
gte-Qwen2-1.5B-instruct (Li et al., 2023b) $\uparrow$ 0.1908 $\uparrow$ 0.1907 $\uparrow$ 0.1794
SFR-Embedding-Mistral (Meng et al., 2024) $\uparrow$ 0.1202 $\uparrow$ 0.1072 $\uparrow$ 0.1181
GritLM-7B (Muennighoff et al., 2024) $\uparrow$ 0.4261 $\uparrow$ 0.4278 $\uparrow$ 0.4155
gte-Qwen2-7B-instruct (Li et al., 2023b) $\uparrow$ 0.1168 $\uparrow$ 0.1288 $\uparrow$ 0.1258
e5-mistral-7b-instruct (Wang et al., 2023b) $\uparrow$ 0.0928 $\uparrow$ 0.1093 $\uparrow$ 0.1145
Linq-Embed-Mistral (Choi et al., 2024) $\uparrow$ 0.0688 $\uparrow$ 0.0649 $\uparrow$ 0.0758
4.4.1. Anchor Strategy: Stock–Stock vs. Theme-Based
In this experiment, we compare two contrastive training configurations: Stock–Stock Alignment (SSA), where both the anchor and the positive are constituent stocks from the same ETF, and the theme-based anchor strategy, where the anchor is a textual description of the theme and the positive is a corresponding constituent stock. While SSA emphasizes local similarity within ETF holdings, the theme-based approach encourages alignment with a more abstract semantic concept.
As shown in Table 3, the theme-based anchor strategy consistently outperforms SSA across all models and evaluation thresholds. For instance, using gte-Qwen2-1.5B, P@3 increases from 0.2216 to 0.4124, and P@5 improves by 0.1907. The improvements are especially notable at lower $k$ values, indicating sharper ranking. These results suggest that using thematic descriptions as anchors offers more effective supervision than intra-stock alignment.
4.4.2. Training Dataset: ETF-only vs. TRS
Separately, we compare training on the Thematic ETF Set, which includes only real ETF data, to training on the more comprehensive TRS. While supervision from ETFs is grounded in actual investment products, it often shows a bias toward popular sectors like IT.
Table 4. Ablation study comparing the Thematic ETF Set (real ETFs only) with our broader TRS dataset. We report the score improvements of our setting over the counterpart for each evaluation metric.
Model P@3 P@5 P@10
bge-small-en-v1.5 (Xiao et al., 2023) $\uparrow$ 0.0705 $\uparrow$ 0.1186 $\uparrow$ 0.1495
bilingual-embedding-large (Conneau et al., 2019) $\uparrow$ 0.0430 $\uparrow$ 0.0485 $\uparrow$ 0.0444
multilingual-e5-large-instruct (Wang et al., 2024) $\uparrow$ 0.0705 $\uparrow$ 0.0578 $\uparrow$ 0.0598
stella_en_1.5B_v5 (Zhang et al., 2025) $\uparrow$ 0.0258 $\uparrow$ 0.0216 $\uparrow$ 0.0207
gte-Qwen2-1.5B-instruct (Li et al., 2023b) $\uparrow$ 0.0086 $\uparrow$ 0.0113 $\uparrow$ 0.0082
SFR-Embedding-Mistral (Meng et al., 2024) $\uparrow$ 0.0515 $\uparrow$ 0.0350 $\uparrow$ 0.0382
GritLM-7B (Muennighoff et al., 2024) $\uparrow$ 0.0327 $\uparrow$ 0.0443 $\uparrow$ 0.0433
gte-Qwen2-7B-instruct (Li et al., 2023b) $\uparrow$ 0.0670 $\uparrow$ 0.0690 $\uparrow$ 0.0479
e5-mistral-7b-instruct (Wang et al., 2023b) $\uparrow$ 0.0498 $\uparrow$ 0.0485 $\uparrow$ 0.0397
Linq-Embed-Mistral (Choi et al., 2024) $\uparrow$ 0.0430 $\uparrow$ 0.0299 $\uparrow$ 0.0294
TRS addresses this by expanding theme coverage through sectoral classification and news-driven augmentation.
As shown in Table 4, models trained on the TRS dataset consistently outperform their ETF-only counterparts across most settings. As an example, bge-small-en-v1.5 gains +0.1186 in P@5, and multilingual-e5-large-instruct improves by +0.0705 in P@3. These findings highlight the importance of broad and balanced thematic coverage in contrastive learning for stock retrieval.
5. Discussion
The results from our experiments demonstrate not only the modeling strengths of THEME, but also its practical value in real-world thematic investing contexts. Rather than focusing solely on academic benchmarks, we highlight how the system’s design and performance translate into tangible benefits for various actors in the investment ecosystem.
First, the retrieval gains shown in Experiment 4.2 support the use of THEME as an idea generation engine. Asset managers can input a high-level thematic description and retrieve a curated list of stocks that extends beyond traditional sector classifications. This facilitates the discovery of emerging or under-recognized companies that may not yet appear in static ETF portfolios. Since our approach consistently outperforms both LLM-based and rule-based retrieval baselines, it offers a reliable starting point for thematic universe construction across discretionary or quantitative workflows.
Second, Experiment 4.3 shows that retrieved stocks are not only thematically relevant but also yield stronger short-term returns when combined with recent performance signals. This confirms the viability of THEME as a component in active investment processes. For example, portfolio managers can use it to periodically update allocations based on a combination of theme relevance and market momentum. Unlike traditional ETF reconstitution, this allows for more timely adjustments while maintaining thematic intent. The integration of semantic and temporal signals makes the system adaptable to both structural and market-driven shifts.
Third, our ablation study in Experiment 4.4 emphasizes two design decisions that directly impact real-world performance: the choice of semantic anchors and the composition of the training dataset. We find that training with abstract theme descriptions as anchors yields better generalization than stock-to-stock alignment. In parallel, training on the expanded Thematic Representation Set (TRS) leads to consistent retrieval gains over using real ETF data alone. This is especially important for practitioners seeking to cover a wide array of investment themes, including those not yet captured in the current ETF landscape. TRS also enables multi-theme associations, ensuring that stocks with diverse business models are not artificially constrained to a single category.
THEME is implemented as a modular, cloud-native platform with API-based access, making it deployable across a range of enterprise environments. Outputs can be integrated into discretionary workflows (e.g., analyst platforms), systematic engines (e.g., ranking or weighting modules), or even client-facing tools for personalized investing. For instance, a retail platform could allow users to enter a theme like ”climate-resilient agriculture” and receive a dynamically updated portfolio reflecting both long-term relevance and short-term return potential.
Looking forward, we are incorporating real-time data sources, such as ESG events, earnings calls, and patent filings, to support additional use cases like intraday thematic scoring, event-driven rebalancing, or early detection of theme dislocation. Rather than replacing existing approaches, THEME serves as a thematic intelligence layer that augments decision-making across the spectrum of passive and active strategies.
Overall, THEME bridges the gap between qualitative narratives and quantitative execution. By translating loosely defined thematic concepts into structured, time-sensitive representations, it enables more adaptive and scalable participation in evolving markets.
6. Conclusion
This work makes three key contributions to thematic investing. First, we propose THEME, a hierarchical contrastive learning framework that integrates semantic and temporal supervision to produce theme-aware stock embeddings. Second, we construct the TRS, a scalable dataset that addresses the coverage limitations of real-world thematic ETFs by incorporating industry classifications and financial news. Third, we demonstrate through extensive experiments that these design choices lead to consistent improvements in retrieval accuracy and portfolio performance across a range of language model backbones, while supporting real-world deployment through modular APIs and customizable interfaces.
THEME addresses a central challenge in thematic investing: finding stocks that are both conceptually aligned and responsive to market conditions. By combining thematic descriptions with recent return patterns, the system generates representations that support both interpretability and quantitative execution.
In our experiments, THEME outperforms extensive baselines including strong LLM and embedding-based baselines in retrieval precision and hit rate (Experiment 4.2), constructs portfolios with improved Sharpe ratios and cumulative returns (Experiment 4.3), and benefits from semantic anchoring and TRS-based training over conventional alternatives (Experiment 4.4). These results confirm that integrating semantic alignment with temporal sensitivity improves the relevance and profitability of thematic stock selection.
Beyond performance, THEME offers practical utility across investment workflows. Its outputs support thematic universe construction, dynamic portfolio rebalancing, and personalized investing. The system’s modular, API-based architecture enables integration into discretionary platforms, systematic engines, or client-facing applications.
While our framework performs well across metrics, several directions remain for future enhancement. Our supervision relies on curated ETF constituents, which may introduce theme bias; incorporating crowd-sourced labels or alternative benchmarks could improve robustness. We also plan to extend the system with richer inputs such as fundamentals, ESG signals, or supply chain structures. Enabling real-time inference based on live news or earnings events would further support intraday applications such as theme-aware risk monitoring or event-driven rebalancing.
Overall, THEME shows how semantic understanding and temporal modeling can be jointly applied to build more adaptive, scalable, and insightful systems for navigating complex investment themes in modern financial markets.
Acknowledgments
This work was supported by National Research Foundation of Korea (NRF) grant (No. NRF-2022R1I1A4069163) and Institute of Information & Communications Technology Planning & Evaluation(IITP) grant (No. RS-2020-II201336, Artificial Intelligence Graduate School Program(UNIST)) funded by the Korea government(MSIT). We would like to thank LG AI Research for providing computational and research support.
GenAI Usage Disclosure
The authors used EXAONE (LG AI Research) solely for language editing purposes, including grammar correction and improvements in readability. No substantive content such as text, data, figures, code, or novel interpretations was generated by the tool. All research ideas, methodology, experiments, analyses, and conclusions are entirely the work of the authors, who bear full responsibility for the accuracy and integrity of the manuscript.